### v1

# Algorithmic resolution of crowd-sourced moderation on X in polarized settings across countries

Abstract
Abstract
Social platforms increasingly transition from expert fact-checking to crowd-sourced moderation, with X pioneering this shift through its Community Notes system, enabling users to collaboratively moderate misleading content. To resolve conflicting moderation, Community Notes learns a latent ideological dimension and selects notes garnering cross-partisan support. As this system, designed for and evaluated in the United States, is now deployed worldwide, we evaluate its operation across diverse polarization contexts. We analyze 1.9 million moderation notes with 135 million ratings from 1.2 million users, cross-referencing ideological scaling data across 13 countries. Our results show X’s Community Notes effectively captures each country’s main polarizing dimension but fails by design to moderate the most polarizing content, posing potential risks to civic discourse and electoral processes.
Keywords: social media, misinformation, moderation, crowd-sourced fact-checking.
Social media platforms have become ubiquitous sources for the consumption of news content  (?), where users may also participate in content creation and dissemination. A persistent challenge in these platforms is the dissemination of misinformation  (?, ?, ?), prompting platforms to navigate a trade-off between freedom of expression on the one hand, and content moderation on the other  (?, ?). While expert fact-checking has been shown to successfully decrease the extent to which people agree with misleading claims  (?, ?, ?), it faces significant limitations in practice: i) professional fact-checking operations cannot scale to meet the volume of content requiring verification  (?, ?), ii) their efforts may be disconnected from online attention  (?), iii) moderation may be perceived as having partisan bias by the public  (?), and iv) they may suffer from low visibility  (?). Additionally, professional moderation may backfire by further “entrenching” polarized individuals  (?). As an alternative, crowd-sourced moderation of content (i.e., systems in which regular platform users signal misinformed or potentially misleading content) has demonstrated that panels of regular users can effectively identify misinformation  (?, ?, ?, ?).
Motivated by these findings, X (formerly Twitter) piloted a crowd-sourced fact-checking system in the United States in early 2021  (?), later deployed worldwide in 2023 under the name Community Notes  (?). This system enables self-selected contributors to attach contextual notes to posts they consider “misinformed or potentially misleading”. Other contributors can then rate these notes based on their subjective assessment of the content and its context. A core challenge for the platform is moderating polarized notes, for which consensus on usefulness may not emerge between raters across the ideological spectrum.
To address this challenge, X first learns to predict how contributors rate notes fitting latent ideology parameters for users and notes, then algorithmically identifies notes that appeal to users across ideological divides rather than only to like-minded individuals  (?, ?). Instead of merely reflecting the idiosyncratic “helpfulness” of a note, contributor ratings are subjective assessments that may incorporate ideological alignment or misalignment between the rater and the claim being discussed. To account for this, X’s Community Notes system infers latent ideological positions for both the notes ($\theta_{n}$) and the raters ($\theta_{r}$) under the key hypothesis that greater ideological alignment increases the probability of a rater rating the note as Helpful.
Computationally, X first converts ratings to numerical values: 1 for Helpful, 0.5 for Somewhat Helpful, and 0 for Not Helpful  (?). Then, to predict these empirical ratings, the system infers several parameters: note latent ideologies ($\theta_{n}$), rater latent ideologies ($\theta_{r}$), individual rater leniency ($\beta_{r}$), the tendency of notes to be rated helpful regardless of raters’ ideologies ($\beta_{n}$), and a global helpfulness baseline ($\beta_{0}$). These parameters are estimated, from the ratings alone, to best predict the rating $\eta$ that rater $r$ will assign to note $n$ through the predictive formula:
$\hat{\eta}_{rn}=\beta_{0}+\beta_{n}+\beta_{r}+\theta_{n}\cdot\theta_{r}.$
Consequently, a note rated Helpful by diverse raters (i.e., with $\theta_{r}$ values spread across positive and negative values) will have a higher $\beta_{n}$ value and lower absolute value $|\theta_{n}|$ when compared to notes rated Helpful by only one side of the spectrum. X’s Community Notes service uses these parameters to algorithmically resolve conflicting ratings by selecting notes with the highest $\beta_{n}$ values  (?) and displays them under the post, visible to all users on the platform  (?).
Results from X’s pilot phase in the United States revealed that Community Notes selected for display garnered broad ideological appeal  (?) and that users exposed to such notes were less likely to share the posts and less inclined to agree with the posts’ misleading claims. Independent research found high agreement between Community Notes and expert fact-checker classifications  (?). Additionally, Allen et al.  (?) showed that political partisanship was highly predictive of contributors’ ratings in the US, with raters preferentially challenging content from those with whom they disagree politically. Following X’s Community Notes global deployment in 2023, subsequent studies corroborated these findings, documenting reduced sharing of flagged content  (?) and accurate sourcing of credible references for notes addressing COVID-19 vaccination misinformation  (?). Moreover, moderation resulting from the Community Notes service was shown to be less prone to backfire effects that entrench individuals  (?, ?) and generated greater user trust than warnings issued by expert fact-checkers  (?).
X’s Community Notes system was developed and tested in the United States  (?, ?), where political disagreement can be effectively modeled using a single ideology dimension  (?, ?). However, the relevance of the internal representation of note and rater ideology —central to the algorithmic moderation process in Community Notes— remains largely unexplored in more general political settings. In countries where political competition is organized along multiple ideological dimensions  (?, ?, ?), what form of polarization is captured by the latent ideology in X’s Community Notes, and how this relates to the performance of this crowd-sourced moderation systems, remains an open question. In this study, we examine X’s Community Notes usage and outcomes across 13 countries and compare the latent ideological space inferred by X’s algorithm with established political dimensions derived from follower networks  (?) and calibrated with political survey data  (?).
Data and definitions
Community Notes data
To enable third-party scrutiny, X publicly discloses all proposed Community Notes, their ratings, the helpfulness status algorithmically assigned to them, and user requests soliciting Community Notes on posts they believe would benefit from additional context.
As of March 1, 2025, X had disclosed data on 1.88 million Community Notes, which received 135.17 million ratings from 1.15 million contributors. Additionally, 1.49 million X users submitted 4.83 million requests for Community Notes associated with 1.91 million posts. Following X’s preprocessing protocol  (?, ?), we retained notes with at least 5 ratings and contributors who rated at least 10 notes. This resulted in 1.45 million notes associated with 886 370 X posts, rated 133.20 million times by 821 250 contributors (see descriptive statistics in SM.SB.1); 59.5% of ratings classify the proposed notes as Helpful, 37.5% as Not Helpful, and 3.0% as Somewhat Helpful. To enable contextual analysis beyond ratings, the posts associated with 83.5% of proposed notes were retrieved, while the remainder had been deleted by March 2025 (see SM.SB.2). Hereafter, we use “notes” to refer to Community Notes flagging the original post as “misinformed or potentially misleading” (73.0% of all notes; the others are not meant for public display but allow contributors to argue why a note is not needed; see SM.SA).
Inferring the parameters of the Community Notes model
We applied the regularized matrix factorization outlined in X’s publication  (?) and implemented in their open-source code  (?), to infer the global helpfulness bias $\beta_{0}$, rater biases $\beta_{r}$, note biases $\beta_{n}$, and the latent ideologies of notes and raters, $\theta_{n}$ and $\theta_{r}$ (see details in SM.SF). The parameters mirror those reported in X’s publication  (?) and align closely with note outcomes publicly displayed on X. Quantitatively, the note biases $\beta_{n}$ we inferred from ratings predict whether a note was algorithmically selected as Helpful by X (hereinafter, achieving Helpful Status) with an AUC of 0.92, and predict Not Helpful Status with an AUC of 0.97 (see the validation of model parameters in SM.SF.3).
Multidimensional ideology scaling of accounts on X
To compare the ideological scale inferred by X’s system with established ideological dimensions across countries, we rely on the method developed by Ramaciotti et al.  (?), which adapts Barberá’s  (?) methodology to multidimensional settings. Specifically, we identify X accounts following Members of Parliament (MPs) from 23 countries across five continents, retaining those that followed at least 3 MPs and were followed by at least 25 accounts, following Barberá’s protocol  (?, ?, ?). We then infer, through Correspondence Analysis, a multidimensional latent space arranging MPs and their followers by homophily (see details in SM.SE.1).
For the analysis of Community Notes latent ideology parameter, we consider the position of X accounts along $\delta_{1}$, the first principal component of the Correspondence Analysis, which best explains which MPs they follow. Additionally, we map the positions of accounts onto interpretable political dimensions of the Global Party Survey  (?) —an expert survey positioning parties along several dimensions endowed with a reference frame making them comparable across countries— namely i) the Left-Right dimension, a leading dimension of polarization relevant for studies of misinformation in social media  (?, ?, ?), and ii) the Anti-Elite dimension, which measures criticism toward elites and institutions and has been proposed as an alternative explanation for the dissemination of misinformation online  (?, ?, ?). When comparing latent ideologies $\theta_{n}$ and $\theta_{r}$ with Left-Right and Anti-Elite dimensions, we focus on 13 countries where at least a thousand X accounts with inferred political positions have authored posts with proposed Community Notes. These countries represent a geographically diverse sample (see SM.SE.3): United States, United Kingdom, Japan, Spain, France, Brazil, Canada, Germany, Argentina, Israel, Australia, Poland, and Mexico.
Results
Notes and ratings
Community Notes usage reveals highly unequal distributions (see SM.SB.1). Among the 1.1 million users self-enrolled in the program as of March 2025, 28.9% have rated fewer than ten notes. Just 1% of the 405 thousand contributors with note-writing privileges authored 31.2% of all proposed Community Notes. Similarly, 35.8% of all proposed Community Notes are associated with posts published by only one thousand X accounts (see most commented accounts per country in SM.SC.5.1). Of the posts for which users requested Community Notes, 79.3% were flagged by only a single user, and a note was proposed for only 12.5% of these posts.
Figure 1.A reveals the worldwide usage of Community Notes. The majority of notes relate to the United States, followed by Japan, the United Kingdom, Brazil, and France (see the country segmentation methodology in SM.SC.5). In the European Union, the number of Community Notes contributors per country correlates strongly with the number of active logged-in users (Pearson $\rho=0.823,~{}p<10^{-4}$), as disclosed under European transparency obligations.
96.2% of Community Notes, after X’s preprocessing  (?, ?), contain URLs pointing to supporting sources that aim to contextualize the original post. To examine news media citation practices in Community Notes, we compiled a list of 2 134 global outlets from BBC Media Guides, supplemented by manual review of the most frequently referenced domains in notes. These established media outlets constitute the predominant sources in notes, appearing in at least 25.9% of notes, followed by X posts (17.4%, where contributors cite other posts as evidence) and Wikipedia articles (8.9% of notes).
Additionally, we curated a list of web domains from professional fact-checking organizations that are International Fact-Checking Network signatories or appear in Duke Reporter’s Lab directory, revealing that expert fact-checking articles appeared in only 3.5% of proposed notes (see the detailed curation of different types of sources in SM.SC.2).
Figure 1: Community Notes usage and outcomes. (A) Fraction of Community Notes proposed per country. To be associated with a country, a note must be predominantly evaluated by contributors who typically rate notes in that country’s language and that reference national websites (see SM.SC.5.1). (B) Cumulative distribution of time needed for a proposed note to reach Helpful Status via X’s algorithm, measured from the publication of the original post and from the publication of the note (truncated tails after 36 hours, representing 9.2% and 17.9% of notes, respectively). (C) Fraction of notes reaching Helpful Status in the US (left) and in the US when referencing expert fact-checks (right), segmented by the ideological leaning of note authors. Error bars show 95% CI via bootstrapping over notes and fact-checking organizations (see SM.SG.1.1).
Political content accounts for 65.2% of posts and 76.6% of Community Notes ratings, with sports as the second most frequent category (6.3% of posts) (see complete distribution and validation in SM.SC.3). To further characterize notes discussing politics, we use a topic model  (?) trained on political party manifestos from the Manifesto Project  (?). Among the topics included in this scheme, the ones most discussed in notes are Political Authority (25.6% of posts, including praise and criticism of specific parties), Law and Order (12.4%, including mention of crimes, the police, and the judiciary), Equality (7.2%, mentioning underprivileged groups, social classes, economic redistribution and discrimination), Environmental Protection (6.6%, including natural resources, natural preservation, and animal rights) and Freedom and Human Rights (6.2%, mentioning freedom of speech, press and assembly) (see a detailed distribution and validation of political topics in SM.SC.4).
Finally, we observe in Figures 2.A & 2.B that, across analyzed countries, Community Notes are requested and proposed under posts authored by accounts throughout the political spectrum. Of the non-deleted posts that received proposed Community Notes and have authors with inferred ideological leanings, 58% are from right-leaning accounts in the United States, 54% in Japan, and 53% in France (see left-right leaning distribution per country in SM.SE.3).
Algorithmic resolution outcomes
We now examine the status assigned by X’s Community Notes algorithm to proposed notes.
Of all proposed Community Notes arguing the original post is “misinformed or potentially misleading”, only 11.97% garner sufficient favorable ratings from “diverse” contributors to attain Helpful Status and be made visible on X. The remaining notes are not publicly displayed, either because consensus determined them to be Not Helpful (2.81%; hereinafter, achieving Not Helpful Status), or because they lacked sufficient ratings from diverse contributors to reach a conclusive status (85.22%). 17.1% of posts for which at least 5 users requested Community Notes ultimately received a proposed note that reached Helpful Status. Between January and March 2025, half of the notes that achieved Helpful Status did so within 4h47min from note submission and 15h17min after the original post publication; see distribution in Figure 1.B.
At the account level, we observe significant variability in the proportion of posts with notes achieving Helpful Status. Politically divisive accounts show fewer notes achieving Helpful Status, for example, those of Elon Musk or Kamala Harris’s campaign have 3% or less than 0.1% of their annotated posts receiving such notes, respectively. Conversely, parody, entertainment, and conspiracy theory accounts exhibit a high proportion of notes achieving Helpful Status (see lists of salient accounts and metrics in SM.SD.3). More generally, 9.7% of annotated posts authored by accounts that held blue check verification prior to X’s 2022 policy change and were deemed of “public interest”  (?) received notes with Helpful Status, compared to 14.0% for unverified accounts; computations restricted to non-deleted posts.
Similarly, we observe that notes discussing politically divisive issues struggle to garner consensus and achieve Helpful Status across each analyzed country, while those reporting scams and other misleading practices commonly do (see SM.SD.2 for country-specific lists of over- and under-represented terms in notes with Helpful Status). For instance, in the United States, only 6.0% of notes discussing presidential candidates and political parties, 5.5% of those addressing the Israel-Palestine conflict, and 11.7% of those covering the Russia-Ukraine war reach Helpful Status. In contrast, 39.0% of notes reporting scams, fraud, and impersonation achieve this status (see SM.SD.2 for a curation of keywords used).
Figure 2: Distribution of Left- and Right-leaning X accounts that authored posts for which (A) Community Notes were requested by users; (B) Community Notes were proposed by contributors; and (C) Community Notes reached Helpful Status. Statistics computed over non-deleted posts authored by accounts with an inferred ideological leaning (see SM.SE.1).
Finally, as reported in Figure 2.C, posts with notes that reached Helpful Status are authored by accounts across the political spectrum in the 13 countries for which we computed ideological positions (see SM.SE.3 for Left-Right distributions across all countries). In the United States, 8.3% of annotated posts authored by Left-leaning users receive notes that achieve Helpful Status, compared to 12.9% for Right-leaning users, as observed in  (?). However, this disparity should not be interpreted as definitive evidence that content from Right-leaning users is flagged more frequently. Since our analysis is restricted to non-deleted posts, the observed difference could be explained, at least partially, by differential deletion patterns, whereby Left-leaning users may delete their posts with Helpful Status notes more frequently than Right-leaning users (see discussion in SM.SG.1.1).
Latent ideological dimension in Community Notes
In the United States, the single-dimensional latent ideological space learned by X’s Community Notes system closely aligns with users’ Left-Right political leanings, notes with $\theta_{n}>0$ appear under posts published by Left-leaning accounts, while notes with $\theta_{n}<0$ are associated with Right-leaning accounts (Fig. 3.A). In particular, 95.0% of notes associated with posts authored by Republican members of Congress have $\theta_{n}<0$, while 94.6% of those associated with Democratic members of Congress have $\theta_{n}>0$ (see the average latent ideologies $\theta_{n}$ of notes associated with key political figures across multiple countries in SM.SG.1). Complementarily, contributors with $\theta_{r}<0$ tend to rate notes as Helpful when they appear under Left-leaning accounts but rate notes as Not Helpful when they appear under Right-leaning accounts (Fig. 3.B), and the opposite pattern holds for raters with $\theta_{r}>0$ (Fig. 3.C), aligning with the observations made by Allen et al. during the pilot program  (?). In the United States, the Left-Right political leaning of users predicts the sign of the latent ideologies $\theta_{n}$ of the notes associated with their posts, with an AUC of $0.822$. Additionally, news outlets cited in Community Notes align ideologically with contributor leanings; we found a Pearson correlation of $\rho=0.744$ ($p<10^{-4}$) between Ad Fontes media bias ratings  (?) and the latent ideology $\theta_{r}$ of contributors citing these media outlets as sources in their notes (see SM.SG.1.1).
In the United States, expert fact-checking articles are used as sources by contributors across the ideological spectrum, though used predominantly ($60.3\%~{}[58.0\%,~{}62.8\%]$) by Left-leaning contributors ($\theta_{r}<0$). As displayed in Figure 1.C, Community Notes incorporating expert fact-checking sources attain Helpful Status at significantly higher rates ($21.3\%~{}[18.3\%,~{}24.1\%]$) than average, with this effect more pronounced for notes authored by contributors deemed Right-leaning according to the Community Notes model ($\theta_{r}>0$), $26.9\%~{}[23.3\%,~{}30.8\%]$ versus $17.6\%~{}[14.0\%,~{}20.7\%]$ for Left-leaning contributors (95% CIs obtained by bootstrapping over notes and fact-checking organization; see SM.SG.1.1).
Figure 3: (A to C) Distribution of Left-Right leaning of X accounts (A) segmented by the sign of the community note latent ideology $\theta_{n}$, (B) rated helpful/not-helpful by raters with negative latent ideology $\theta_{r}<0$, (C) rated helpful/not-helpful by raters with positive latent ideology $\theta_{r}>0$. (D to G) X accounts positioned in the (Left-Right; Anti-Elite) 2D plane, segmented by the majority sign of the latent ideology $\theta_{n}$ of their associated Community Notes (red for $\theta_{n}<0$ and blue for $\theta_{n}>0$). The direction that best separates notes by majority outcome, learned through logistic regression, is displayed as $CN$ along with the corresponding AUC score for sign prediction based on account positions in the 2D plane (standard deviation from 10-fold cross-validation). The direction structuring each country’s political landscape $\delta_{1}$ is projected onto the plane.
The latent ideology computed by the Community Notes model ($\theta_{n}$ and $\theta_{r}$) shows varying degrees of alignment with the Left-Right and the Anti-Elite dimensions computed with ideology scaling and political surveys. Figures 3.D to 3.G map X users with proposed Community Notes along the Left-Right and Anti-Elite dimensions in the United Kingdom, Japan, France, and Brazil. The latent dimension learned by X’s system ($CN$ in Fig. 3) —along which notes and raters are positioned— aligns closely with each country’s most structuring direction ($\delta_{1}$ in Fig. 3) in this Left-Right/Anti-Elite plane. We quantify this alignment by measuring how well the position of a user along $\delta_{1}$ predicts the sign of the latent ideology $\theta_{n}$ of notes associated with their posts. AUC values range from 0.850 in Poland to 0.729 in Israel, averaging $0.808\pm 0.037$ across 13 countries. Complementarily, user positions in the Left-Right/Anti-Elite plane predict the sign of the latent ideology $\theta_{n}$ of notes with an AUC of $0.813\pm 0.035$ across all 13 countries (see breakdown for all countries in our study in SM.SG.2).
We find that the direction learned by X’s system to structure content veracity disputes ($CN$ in Fig. 3) aligns with the Left-Right dimension but cannot always be reduced to it, as anticipated by recent literature  (?, ?). Instead, it may present additional alignments with attitudes towards elites and institutions (and, to a lesser degree, possibly additional issue dimensions not considered in our study), as highlighted by the cases of Japan and Brazil (Fig. 3.F & 3.G).
Discussion
We combined X’s Community Notes data, the algorithmic representation learned by this service, and multi-country ideological scaling for characterizing its operation in polarized settings. We showed, in 13 countries, that notes are requested, proposed, and displayed under posts authored by accounts throughout the political spectrum. The majority of proposed Community Notes discuss politics, and predominantly rely on news outlet articles as sources, followed by X posts and Wikipedia articles.
X’s Community Notes system positions notes and raters along a learned latent ideological axis that we show closely aligns with each country’s most structuring political dimension. This latent axis corresponds to the Left-Right dimension and, with differences across countries, to an Anti-Elite dimension. These findings provide the first empirical support for the key hypothesis underlying X’s Community Notes’ design in a global comparative context.
Our results highlight shortcomings of this crowd-sourced system. While notes displayed on X (i.e., having reached Helpful status) have been shown to align with professional fact-checking assessments  (?, ?), factual but polarizing content may still fail to be resolved as Helpful. Only one-tenth of proposed notes ultimately appear on X, with even lower rates for notes associated with politically prominent figures, popular accounts, and divisive issues. More generally, and by design, the most polarizing content relating to the structuring dimension of national politics typically cannot be algorithmically resolved by Community Notes. This limitation is especially relevant to events that activate such structural lines of division, such as national elections, pointing toward the potential complementarity between expert and crowd-sourced fact-checking, as originally intended by X  (?).
Furthermore, the fraction of polarizing content that will remain unresolved (regardless of factuality) is a design choice made by the platform, both explicitly by fixing the ideological diversity threshold for a note to reach Helpful Status and implicitly by adjusting the regularization strength of the various latent parameters of the Community Notes model.
We acknowledge limitations in our study. In particular, our analyses rely on a snapshot of the data produced by the service until March 1, 2025, and disregard potential fine-grained temporal dynamics. Additionally, our segmentation of notes and raters per country is based on their involvement in the national political discussions by selecting those that follow MPs, combined with their use of the country’s national language. Given the absence of additional metadata disclosed by the platform, we argue that this heuristic is reasonable and necessary for analyzing ideological dynamics in Community Notes.
Importantly, our post analysis was limited to content that remained available at the time of collection in March 2025, introducing a potential bias in the analysis, as posts having notes achieving Helpful Status are more likely to be removed by their authors  (?). Finally, X’s Community Notes contributors are self-selected, and the platform provides no information to characterize this population demographically, let alone politically.
As platforms transition from expert fact-checking to crowd-sourced moderation, with YouTube  (?), TikTok  (?), and Meta  (?) deploying similar systems —the latter explicitly using X’s open source algorithm as the basis of their rating system— this study provides a timely and comparative characterization of their operation in polarized settings across diverse countries. A number of arguments exist for considering alternative to expert fact-checking, as these approaches accumulate challenges with regard to the required scale, reactivity, and efficiency  (?), but also concerning perceived legitimacy and credibility  (?). Yet, our results show that X’s Community Notes system, by relying on consensus rather than factuality, presents inherent limitations in addressing polarizing content. The use of crowd-sourced moderation as a substitute for expert fact-checking therefore warrants careful consideration and calls for risk assessment comparing the moderation outcomes of these approaches over contested polarizing content widely disseminated during events such as elections.
Acknowledgments
Funding
This work has been partially funded by the “European Polarisation Observatory” (EPO) of CIVICA Research (co-)funded by EU’s Horizon 2020 programme under grant agreement No 101017201, by European Union Horizon program project “Social Media for Democracy” under grant agreement No 101094752 (www.some4dem.eu), by the Very Large Research Infrastructure (TGIR) Huma-Num of CNRS, Aix-Marseille Université and Campus Condorcet, and by Project Liberty Institute project “AI-Political Machines” (AIPM). P.B. acknowledges support from the Jean-Pierre Aguilar fellowship from the CFM Foundation for Research.
Ethics and compliance declarations
Our study did not involve experimentation with human subjects. All data used used in our study is publicly available through X’s API and through the Community Notes program at https://x.com/i/communitynotes/download-data.
The processing of political attributes data was declared on 19 March 2020 and 15 July 2021 at the registry of data processing at the Fondation Nationale de Sciences Politiques (Sciences Po) in accordance with General Data Protection Regulation 2016/679 (GDPR) and X policy. For further details and the respective legal notice, please visit the web page of project EPO: medialab.sciencespo.fr/en/activities/epo/. The processing of the data has been approved by the Research Ethics Committee of the Paris Institute of Political Studies in its decision nº2023-038.
Supplementary Materials for
Algorithmic resolution of crowd-sourced moderation on X in polarized settings across countries
Paul Bouchaud, Pedro Ramaciotti
[…]
Appendix A The Community Notes service
The declared aim of X’s Community Notes is to improve information quality by “empowering people on X to collaboratively add helpful notes to posts that might be misleading”  (?). The system combines collaborative annotation with an algorithmic selection process that determines which notes are displayed publicly alongside the posts they moderate.
Figure S1: (A) An example of a Community Note flagging a post as “misinformed or potentially misleading”, prompting contributors to rate its “helpfulness”. (B) An example of a Community Note that has reached Helpful Status and is displayed beneath the original post on X.
X users, having signed up as contributors, can write notes explaining why a post may be “misinformed or potentially misleading”. Other contributors then rate these proposed notes as Helpful, Somewhat Helpful, Not Helpful; as displayed in Figure S1.A. If sufficient contributors from “diverse viewpoints” rate a note as Helpful, X assigns it the status of being Helpful (denoted as Helpful Status through this article to differentiate from a rating given by a user to a note) and it becomes visible to all users on X; as displayed in Figure S1.B. Conversely, if enough contributors from different perspectives rate a note as Not Helpful, the note remains hidden from public view. Contributors who repeatedly author notes assigned Not Helpful Status may lose their ability to contribute further notes. Notes that have not yet received sufficient ratings from contributors with diverse perspectives remain invisible to users outside the Community Notes program. This ensures that only well-vetted content reaches the broader X audience.
Figure S2: (A) An example of a Community Note arguing that the original post is not misinformed or potentially misleading” and that no notes are needed. (B) Signal displayed under a post for which regular X users requested a community note to be proposed. (C) Taxonomy of outcomes for Community Notes; only notes arguing that the original post is misinformed or potentially misleading” and that reach Helpful Status are displayed on X.
The system also includes a type of note called “Note Not Needed”, which allow contributors to argue against the necessity of a proposed note; as displayed in Figure S2.A. These notes are never displayed publicly on X regardless of their status, serving solely as internal discussions among contributors about whether a note addressing alleged misinformation is warranted.
Since June 2024, X has expanded participation by allowing everyone on the platform with a verified phone number (and not only users enrolled in the Community Notes program) to request Community Notes on posts they believe would benefit from additional context  (?). When sufficient requests accumulate, “Top Writers” (contributors recognized for writing a significant number of notes that are found Helpful by others) receive alerts and can choose to propose notes; as displayed in Figure S2.B. This mechanism enables broader user participation while letting contributors know where notes might be most helpful.
Appendix B Data collection
B.1 Community Notes
B.1.1 Notes and ratings
X publicly and continuously releases all Community Notes contributions enabling third-party scrutiny  (?). We consider in this study 1.88 millions notes submitted prior to March 1 2025, and their associated 135.17 million ratings. We display in Figure S3, the number of ratings submitted daily between November 2022 —when the program was renamed Community Notes— to March 1, 2025. The number of ratings is observed to increase within this period, as the program was deployed in more countries  (?).
Figure S3: Number of ratings submitted daily on X’s Community Notes between November 2022 and March 2025.
Figure S4 shows the distribution of time spans between first and last ratings per rater, along with average weekly rating frequency; restricted to raters with multiple submissions. Results demonstrate sustained engagement: 82.6% of raters maintained activity across at least six months, while 40.2% averaged more than one rating per week.
Figure S4: Distribution of time spans between first and last ratings per rater (left), and of average weekly rating frequency (right); restricted to raters with multiple submissions
Up to March 2025, X’s Community Notes were rated by 1,150,257 raters, with a median of 29 (mean: 117.5) notes rated per rater (where a rater is defined as a contributor who rated at least one Community Note). Overall, 28.6% of raters evaluated fewer than 10 notes. The cumulative distribution of ratings per rater is displayed in Figure S5.B.
At the note level, Community Notes received a median of 26 ratings (mean: 77.2), with 16.9% receiving fewer than 5 ratings, we display the cumulative distribution of rating per note on Figure S5.C.
To unlock note writing privileges, a contributor must achieve a “rating impact” of at least 5, meaning they must rate 5 notes as Helpful (or Not Helpful) that eventually reach Helpful Status (or Not Helpful Status). As of March 2025, 36.0% of users enrolled in the Community Notes program had note writing privileges; corresponding to 405k contributors. Among those, 253,002 authored at least one Community Notes, with a median of 3 (mean: 7.4) notes per author (where an author is defined as a contributor who submitted at least one note). Figure S5.A shows the cumulative distribution of notes per author. The distribution exhibits high inequality, with the top 10% of authors writing 57.6% of all Community Notes.
Figure S5: Cumulative distribution of notes per author (A), rated notes per rater (B) and ratings per note (C).
Following X’s pre-processing procedure  (?, ?), we retained raters who rated at least 10 notes and notes with at least 5 ratings; resulting in a final dataset of 1.45 million notes rated 133.2 million times by 821,251 contributors. We will consider this set of notes and ratings when training the Community Notes algorithm in Section F.
B.1.2 Note requests
As of March 1, 2025, 1.49 million X users have submitted 4.83 million requests for Community Notes across 1.91 million posts. Among users who have requested notes at least once, 65.0% submit only a single request; see cumulative distribution in Figure S6.A. Similarly, 79.3% of flagged posts receive only one request for a note, while 93.5% receive fewer than 5 requests; see cumulative distribution in Figure S6.B. We display in Figure S6.C the fraction of posts receiving a proposed Community Notes in function of the number of requests they received; computation restricted to posts having receive request(s).
Figure S6: Cumulative distribution of notes requested per author (A), and number of requests per post (B). Fraction of posts receiving a Community Note as a function of the number of requests they received; computation restricted to posts having received requests; truncated tail (above 50 requests) represents 0.4% of requested posts (C).
B.2 Posts associated with notes
To comply with platform policies, user privacy rights, and deletion requests, X’s Community Notes dataset only discloses post IDs associated with notes, without further information on post content or author. To enable contextual analysis beyond ratings, we collected post content and author information, associated to those notes. As of March 2025, 616,104 X posts associated with 83.5% of non-deleted Community Notes were collected. The remaining posts were unavailable at the time of collection, i.e., having been removed.
Additionally, out of 966k posts with at least two user requests for a Community Note, we collected 804k posts (83.2%); the remaining posts were unavailable at the time of collection in May 2025.
Importantly, it should be noted that such collection process (as opposed to live stream collection or access to data reserved to the platform) introduces an inherent observational bias. Indeed, some analyses are limited to posts that remain available at collection time, which then captures the tendency of users to remove their posts after the note reached Helpful Status. Specifically, we observe that 33.4% of posts with notes having reached Helpful Status were deleted by the time of collection, compared to only 15.1% for posts associated with note with Needs More Ratings Status or Not Helpful Status. This observational bias should be considered when interpreting the following results. Analyses that rely solely on the content of Community Notes, however, such as source analysis, are not marred by such limitation.
Figure S7: Cumulative distribution of number, per X account, of posts with requests for Community Notes (A) and with a submitted Community Notes (B); conditional on having at least one request/annotated post. (C) Cumulative distribution of annotated posts’ number of likes and reposts.
The 616,943 collected posts associated with a proposed Community Notes (hereinafter refer to as “annotated posts”) were authored by 140,913 X accounts, averaging 7.8 notes per author (with median equal to 2); among accounts with at least one annotated post. Elon Musk’s account (@elonmusk) had the highest number of notes, with over nineteen thousand notes proposed by the community under more than 3,300 posts, followed by conspiracy theorist account
@BGatesIsaPyscho with over four thousand notes proposed by the community under more than 2,700 posts; we display these account profiles in Figure S8. Similarly, Elon Musk’s account (@elonmusk) had the highest number of posts (at least 3,087) with Community Notes being requested, followed by Kamala Harris’s campaign account (@KamalaHQ) with at least 1,203 posts with requests. We display in Figure S7.A the distribution per account of posts with requests for Community Notes.
(a) Profile of X account: @elonmusk.
(b) Profile of X account: @BGatesIsaPyscho
(c) Profile of X account: @KamalaHQ
Figure S8: Profiles of the X accounts having received the highest number of Community Notes, and number of requested Community Notes as of March 2025.
Figure S7 shows the distribution of number of annotated post per X account, along with the distributions of number of likes and reposts for annotated posts. The collected posts averaged 13,693 likes and 2,224 reposts (medians of 2,705 and 480 respectively). We found a positive Spearman correlation, $\rho=0.444,p<10^{-4}$, between a post’s number of likes and the number of ratings of notes associated to it; and similarly with the number of reposts $\rho=0.437,p<10^{-4}$.
Appendix C Characterization of notes
C.1 Misleading nature of posts
Upon submitting Community Notes, contributors respond to “Why do you believe this post may be misleading?”. For each (non-exclusive) proposed reason, we report the corresponding fraction of notes and the proportion of them reaching Helpful Status:
  • “It is a misrepresentation or missing important context”: 61.7% of notes; 11.5% reached Helpful Status
  • “It contains a factual error”: 56.4% of notes; 12.1% reached Helpful Status
  • “It presents an unverified claim as a fact”: 35.5% of notes; 10.7% reached Helpful Status
  • “It contains outdated information that may be misleading”: 20.8% of notes; 12.0% reached Helpful Status
  • “Other”: 13.4% of notes; 12.7% reached Helpful Status
  • “It contains a digitally altered photo or video”: 9.2% of notes; 17.4% reached Helpful Status
  • “It is a joke or satire that might be misinterpreted as a fact”: 7.2% of notes; 13.4% reached Helpful Status
Statistics were computed across the whole set of Community Notes arguing the original post is “misinformed or potentially misleading”.
C.2 Web sources used in notes
Community Notes contributors are encouraged to use sources to justify their proposed notes  (?). To characterize their usage, we extract all URLs used in Community Notes qualifying the original post as “misinformed or potentially misleading”; resolving X’s link shortener (t.co) and Google “amp” redirections. Overall, 96.2% of Community Notes contain URLs pointing to supporting sources that aim to contextualize the original post. We identified 76.0 thousand unique domains; Figure S9.B presents the distribution of Community Notes referencing them.
As displayed in Figure S9.A, it emerges that x.com (aggregated with twitter.com) is the domain most frequently referred to, followed by wikipedia.org, youtube.com (aggregated with youtu.be) and bbc.com (aggregated with bbc.co.uk) and reuters.com.
References to x.com can be further disaggregated into user-generated content (posts, profiles, search results) and platform documentation. Specifically, 18.3% of Community Notes referencing to x.com or twitter.com refer to one of the following ressources: communitynotes.x.com,
help.x.com, blog.x.com, business.x.com, and
x.com/en/tos (combined with twitter.com variations).
Figure S9: (A) Cumulative distribution of number of notes referencing each domains. (A) Number of notes referencing the 10 most frequently cited domains. Statistics computed on the set of Community Notes processed by X’s Community Notes algorithms, following their pre-processing procedure  (?, ?).
C.2.1 News outlets
We manually inspected the first 500 web domains most cited in notes to identify those pointing towards media outlets, excluding website publishing exclusively fact-checks, and report on the first 139 most cited among them:
reuters.com, apnews.com, news.yahoo.com, news.yahoo.co.jp, bbc.co.uk, nhk.or.jp, bbc.com, theguardian.com, cnn.com, nytimes.com, impress.co.jp, forbes.com, globo.com, vice.com, usatoday.com, washingtonpost.com, nbcnews.com, ft.com, dailymail.co.uk, npr.org, news.yahoo.com, timesofisrael.com,
newsweek.com, uol.com.br, cbsnews.com, afp.com, asahi.com, aljazeera.com, independent.co.uk, abcnews.go.com, infobae.com, lemonde.fr, politico.com, telegraph.co.uk, foxnews.com, wsj.com, cnbc.com, pbs.org, cbc.ca, france24.com, jpost.com, time.com, francetvinfo.fr, thehill.com, businessinsider.com, elpais.com, cnnbrasil.com.br, liberation.fr, indiatimes.com, euronews.com, elmundo.es, haaretz.com, politico.eu, estadao.com.br, leparisien.fr, bloomberg.com, lefigaro.fr, bfmtv.com, ctvnews.ca, dw.com, eldiario.es, gazetadopovo.com.br, hindustantimes.com, indiatoday.in, jiji.com, lanacion.com.ar, latimes.com, mainichi.jp, metropoles.com, ndtv.com, nikkei.com, nypost.com, spiegel.de, poder360.com.br, thehindu.com, yomiuri.co.jp, ynet.co.il, tagesschau.de, sankei.com, 20minutes.fr, 20minutos.es, aa.com.tr, abc.es, bild.de, br.de, cnn.co.jp, economist.com, elconfidencial.com, haaretz.co.il, i24news.tv, irishtimes.com, lepoint.fr, kyivindependent.com, lavanguardia.com, nationalpost.com, ndr.de, theatlantic.com, theverge.com, zeit.de, wired.com, express.co.uk, vox.com, rtve.es, rte.ie, rp.pl, publico.es,
news.sky.com, itmedia.co.jp, axios.com, aap.com.au,
veja.abril.com.br, economictimes.com, indianexpress.com, standard.co.uk, elespanol.com, clarin.com, tokyo-np.co.jp, ouest-france.fr, welt.de, voanews.com, washingtonexaminer.com, tribune.com.pk, sueddeutsche.de, faz.net, thetimes.co.uk,
dawn.com, tagesspiegel.de, news.com.au, pagina12.com.ar, elperiodico.com, israelhayom.co.il, tf1info.fr, globalnews.ca, radiofrance.fr, timesnownews.com, tvn24.pl, lasexta.com,
msnbc.com, usnews.com
These domains were cited in 220k unique notes (after pre-processing), representing 23.2% of the notes.
To test the exhaustiveness of this list of 139 media outlets we obtained the list of media outlets active around the world, curated by the BBC Media Guides. This list includes 2061 web domains of news outlets, covering 154 countries. Combined with our manually curated list, we establish a set of 2134 unique news domains, of which 979 are used at least once in Community Notes as sources. The addition of this second list of medias did not significantly altered the percentage of notes that used news media articles as sources, when considered the union of both lists. Overall, domains from this extended list of news media outlets are used in 246k Community Notes (after pre-processing), constituting the most common reference type (25.9%) ahead of user-content on X (17.4%) and Wikipedia articles (8.9%), demonstrating the prevalence of editorial journalism in X’s crowd-sourced fact-checking system.
C.2.2 Expert fact-checking
To contrast Community Notes with expert fact-checking, we identified notes citing expert fact-checking sources. To this end, we compiled a list of fact-checking sources by combining Duke University’s Reporters’ Lab directory  (?), which catalogs over 400 geographically diverse fact-checking websites, with news agencies and signatories of the International Fact-Checking Network  (?).
To avoid including URLs of general news coverage sources, we manually curated the most frequently used domains from this compiled list. Specifically, we segmented websites into those that published exclusively fact-checks or maintained dedicated fact-checking sections (e.g., reuters.com/fact-check), and those publishing both general news coverage and fact-checking articles. For news organizations that publish fact-checks alongside general coverage without dedicated fact-checking sections in their URL structure, we required the presence of the regular expression “fact.*check” in URLs; including translations. This filtering approach ensures that, e.g., cnn.com articles are not considered fact-checks except for those matching this pattern, such as, edition.cnn.com/2025/04/23/politics/price-of-eggs-gas-trump-fact-check.
From this curated list of over 500 geographically diverse fact-checking organizations, we identified 46,869 Community Notes referencing them as sources within the entire set of Community Notes published prior to March 1, 2025 (including “Note Not Needed” notes). The most frequently cited fact-checks are published by snopes.com, reuters.com, politifact.com, apnews.com, and usatoday.com.
After following X’s pre-processing guidelines  (?, ?) and excluding “Note Not Needed” notes, Community Notes citing fact-checking websites constitute 3.5% (33,329) of notes considered by the Community Notes algorithm.
C.3 Topics of posts and notes
To characterize the textual content of the proposed Community Notes and their associated posts, we first categorized posts into 19 distinct topics using Antypas et al.’s model  (?), after translating all posts to English using Facebook’s M2M100 translation model  (?). These models were selected for their open-weight architectures and top-tier performance; we refer to  (?) and  (?) respectively for details on model training.
To validate topic interpretation, we present below, for the ten most frequent topics, the 15 words the most over-represented, per the chi-square statistic, in the posts of each topic, compared to the overall corpus.
  • News And Social Concern: trump, israel, president, government, biden, hamas, ukraine, election, israeli, minister, party, police, russia, donald, kamala;
  • Sports: football, league, messi, player, olympics, ronaldo, players, olympic, goals, season, match, cristiano, champions, madrid, sports;
  • Diaries And Daily Life: going, thing, sorry, happy, think, little, woman, looking, thank, friend, birthday, girlfriend, morning, father, child;
  • Other Hobbies: issue, previous, paypay, featured, prophet, trusted, allah, going, outer, sunshine, follow, reply, gujarat, spirit, likes;
  • Science And Technology: iphone, apple, earth, space, solar, technology, android, google, spacex, science, battery, phone, microsoft, tesla, software;
  • Fitness And Health: cancer, vaccine, autism, vaccines, disease, health, vaccination, covid, doctor, medical, vaccinated, study, doctors, brain, effects;
  • Celebrity And Pop Culture: taylor, nicki, swift, drake, diddy, minaj, megan, selena, bieber, meghan, kendrick, ariana, cardi, kanye, rihanna;
  • Business And Entrepreneurs: money, bitcoin, paypay, tesla, company, economy, business, companies, crypto, stock, select, trading, market, million, transfer;
  • Film, Tv And Video: movie, anime, disney, movies, netflix, marvel, video, spider, episode, simpsons, animation, films, superman, series, cinema;
  • Gaming: playstation, games, fortnite, nintendo, gaming, rockstar, console, trailer, roblox, steam, minecraft, gamers, switch, gtavi, starfield.
For brevity, we refer to posts classified as “News and Social Concern” —defined by  (?) as content related to awareness, activism, dialogue, and discussion of social issues and newsworthy events— as “political” posts.
Figure S10 shows the topic distribution of annotated posts (regardless of note status). Political content dominates with $65.2\%$ of posts and 76.6% of ratings. Sports ranks second at $6.3\%$ of posts, followed by “Diaries & Daily Life” (slice of life, everyday content that illustrates personal opinions, feelings, occasions, and lifestyles).
Figure S10: Distribution of topics discussed in annotated posts according to dimosthenis: (left) fraction of posts and (right) fraction of community note ratings associated with the posts. The 10 most frequent topics are displayed; less frequent topics are aggregated as “Other Topics”.
Next, we investigate the relationship between post topics and Community Notes rating outcomes. Figure S11 displays the distribution of note status across topics. Since only one note can be displayed under a post, we prioritize these statuses in the following order: Helpful Status, Not Helpful Status, and Needs More Ratings Status. For instance, if a post has two associated notes one with Helpful Status and the other with Needs More Ratings Status the post is classified as having a note with Helpful Status.
While only 11.9% of annotated political posts, having been retrieved, receive a note with Helpful Status, this rate increases to 18.5% for non-political posts; statistical significance confirmed through one-vs-rest chi-squared tests (p-value $<10^{-4}$).
This observation may support different hypotheses. It may reflect less consensus on political posts compared to other topics. Alternatively, it may reveal a higher propensity of users to delete their posts upon receiving a note with Helpful Status when discussing political issues rather than sports or sciences. Finally, this result may be a direct consequence of X’s Community Notes algorithm. Indeed, as the vast majority of Community Notes are associated with political posts, rater latent ideology $\theta_{r}$ learned by X’s systems is likely to capture their political stance, rather than their views on other topics, such as their favorite soccer team. Consequently, when rating notes on non-political topics, users with diverse “political viewpoints” —but potentially similar views on the specific topic being discussed— may rate the note as Helpful, leading to its selection by X’s Community Notes system. These non-exclusive factors that may explain the lower rate of notes with Helpful Status notes on political posts cannot be definitively tested due to lack of access to deleted posts.
Figure S11: Distribution of note statuses across post topics.
C.4 Classifying political notes
To achieve a finer granularity in our analysis of political posts, we leveraged the Manifesto Project coding scheme  (?) and a language model trained on this taxonomy. Specifically, we employed Manifestoberta  (?), a language model based on the XLM-RoBERTa architecture  (?), trained on 1.6 million annotated statements from the Manifesto Corpus  (?), which provides robust performance in classifying political texts  (?). By translating all political posts to English, through Facebook’s M2M100 translation model  (?), we ensure equivalent classification performance across post original languages.
Manifestoberta provides, for each analyzed document (notes in our case), a probability score of mentioning a topic category among those existing in its topic ontology. We implemented a classification threshold at 0.5, assigning 230k political posts to a Manifesto Project category. To validate the inferred categories, we extracted, for the ten most frequent categories, the 15 most over-represented words, per the chi-square statistic, in the post of each category, compared to the overall corpus. The Manifestoberta topic scheme differentiates topics that address issues with positive or negative attitudes towards them.
  • 305 - Political Authority. trump, president, biden, donald, party, harris, kamala, minister, milei, trudeau;
  • 605 - Law And Order: Positive. police, crime, arrested, violence, murder, prison, shooting, criminal, sexual, officers;
  • 503 - Equality: Positive. trans, black, women, racist, racism, white, gender, transgender, lgbtq, woman;
  • 501 - Environmental Protection: Positive. climate, water, warming, emissions, change, polluted, ocean, animals, pollution, weather;
  • 201 - Freedom And Human Rights. freedom, speech, palestinian, rights, censorship, obtain, israeli, journalists, expression, journalist;
  • 304 - Political Corruption. corruption, corrupt, fraud, money, million, hunter, epstein, scandal, biden, election;
  • 504 - Welfare State Expansion. vaccine, covid, vaccines, vaccinated, deaths, vaccination, cancer, health, colona, pension;
  • 603 - Traditional Morality: Positive. abortion, jesus, church, christian, christians, catholic, christ, christianity, religious, abortions;
  • 202 - Democracy. democracy, elections, parliamentary, voting, votes, election, democratic, electoral, ballots, voters;
  • 104 - Military: Positive. missile, russian, missiles, military, aircraft, hezbollah, ukrainian, fighter, forces, defense.
Figure S12 displays the distribution of Manifestoberta categories in retrieved political posts. Within our corpus, “Political authority”, relating to governance competence and political power (with posts featuring political figures’ names) emerges as the most frequent category with 25.5% of posts and 26.5% of ratings. “Law and order” ranks second, defined as “favorable mentions of strict law enforcement, and tougher actions against domestic crime”, as corroborated by the extracted keywords. “Freedom and human rights” ranks third, which in our corpus of annotated posts primarily addresses free speech censorship, and the Israel-Gaza conflict, as evidenced by the extracted keywords.
Figure S12: Distribution of topics discussed in political posts according to Manifestoberta: (left) fraction of posts and (right) fraction of community note ratings associated with the posts. The 10 most frequent topics are displayed; less frequent topics are aggregated as “Other Categories”.
Figure S13 presents the distribution of notes status associated to political posts across categories. Posts discussing “Freedom and Human Rights”, “Equality”, and “Political Corruption” exhibit the lowest proportion of note with Helpful Status. Conversely, posts about “Environmental Protection”, “Military” and “Law and Order” demonstrate higher consensus, resulting in a greater proportion of notes achieving Helpful Status in X’s Community Notes system.
Figure S13: Distribution of note statuses across category of political posts.
Interpretations of these results are subject to the same caveats outlined in subsection C.3.
C.5 Country segmentation
C.5.1 Attributing notes and raters to countries
To enable an analysis disaggregated by national setting, we segment Community Notes by country. In the absence of explicit metadata from X, we develop a multi-stage approach relying on note language, source websites, and contributor rating patterns.
First, we combined notes language and the top-level domain of websites cited in notes (when present); we display in Figure S14 the distribution of these features. A note is associated with a specific country when it meets two criteria: (1) it is written in one of the country’s national languages, and (2) it cites either a national news outlet or a website under the national top-level domain. For instance, a note written in Japanese and containing a link to sankei.com (a Japanese newspaper) or to any website with the .jp domain extension (e.g., nhk.or.jp) will be assigned to Japan. From the complete set of Community Notes published prior to March 2025, this first conservative heuristic leave 71.2% of notes unassigned —those with missing or multiple inconsistent national identifiers, e.g., notes solely using x.com as source.
Figure S14: Distribution of Community Notes language (A) and top-level domains of websites used as sources (B). Statistics compiled from 1.2 million Community Notes submitted prior to March 1, 2025, considering the original post as “misinformed or potentially misleading”.
Second, we map raters to countries based on their rating behavior. Each rater is assigned to the country whose notes (as identified in the first pass) they rate most frequently, restricting our analysis to rater who have rated at least 5 notes assigned to a country in the first pass.
Finally, we extend this labeling to all notes and raters through an iterative classification. Unassigned notes receive country labels based on the majority country of their raters, while raters are assigned to countries based on which country’s notes they most frequently evaluate.
This multi-stage approach is necessary to expand the initial conservative but robust country assignments (based on explicit website and language indicators) to the complete dataset using rating patterns. Ultimately, a note is associated with a country when it is rated predominantly by contributors who typically rate notes with that country’s language and national websites. Figures 1 and S15 display the number of Community Notes written and the number of contributors across countries.
Figure S15: Distribution of estimated number of Community Notes contributors per country. See C.5.1 for methodology details.
There are inherent limitations to this approach, relying on the assumption that raters predominantly evaluate content from their own countries and fail to capture cross-country engagement of some raters. Yet, in the European Union, we observe a strong Pearson correlation ($\rho=0.823,~{}p<10^{-4}$) between the number of Community Note contributors per country and the number of active logged-in users, as disclosed under European transparency obligations.
C.5.2 Accounts receiving most notes per country
To corroborate the country segmentation, we report in ten countries, the X accounts having the highest number of submitted Community Notes (irrespective of their “helpfulness” status):
  • United States: @elonmusk, @BGatesIsaPyscho,
    @jacksonhinklle, @iluminatibot, @EndWokeness, @KamalaHQ, @JoeBiden, @VigilantFox, @libsoftiktok, @POTUS
  • Japan: @kharaguchi, @kirinjisinken, @tweetsoku1, @jhmdrei, @himuro398, @JINKOUZOUKA_jp, @mattariver1, @livedoornews, @newssharing1, @CS60osaka1
  • United Kingdom: @Conservatives, @RishiSunak, @BladeoftheS, @PolitlcsUK, @Nigel_Farage,
    @BGatesIsaPyscho, @darrengrimes_, @Keir_Starmer, @TRobinsonNewEra, @elonmusk
  • Spain: @vitoquiles, @el_pais, @CapitanBitcoin, @eldiarioes, @elmundoes, @FonsiLoaiza,
    @sanchezcastejon, @wallstwolverine, @abc_es, @AlertaNews24
  • Brazil: @choquei, @folha, @LulaOficial, @Metropoles, @GloboNews, @felipeneto, @siteptbr, @gleisi, @CR7Brasil, @NewsLiberdade
  • France: @CerfiaFR, @silvano_trotta, @BFMTV,
    @JLMelenchon, @AlertesInfos, @RimaHas, @ALeaument, @f_philippot, @sandrousseau, @franceinfo
  • Germany: @jreichelt, @tagesschau, @Markus_Soeder, @niusde_, @ZDFheute, @Karl_Lauterbach, @Alice_Weidel, @polenz_r, @SHomburg, @Bundeskanzler
  • India: @MrSinha_, @zoo_bear, @INCIndia,
    @IndianTechGuide, @dhruv_rathee, @ANI, @rishibagree, @INCKerala, @RahulGandhi, @TimesAlgebraIND
  • Argentina: @JMilei, @porqueTTarg, @therealbuni, @ElTrumpista, @laderechadiario, @infobae,
    @clarincom, @TugoNews, @GordoDan_, @TraductorTeAma
  • Israel: @YinonMagal, @netanyahu, @bezalelsm,
    @itamarbengvir, @ShaykhSulaiman, @jacksonhinklle, @Now14Israel, @TallyGotliv, @rothmar, @amit_segal,
Manual inspection confirms that these accounts align with their inferred countries of origin. Notably, Elon Musk’s account @elonmusk ranks among the most commented in both the United States and the United Kingdom. Notes associated with Elon Musk’s posts classified as United Kingdom are attached to content discussing UK politics and contain references to UK news outlets, while notes associated with US-related content are classified accordingly. Figure S16 illustrates examples of posts whose associated notes were classified as United Kingdom. Similar patterns emerge with @BGatesIsaPyscho, an account that publishes both US and UK related content, and @jacksonhinklle, whose content addresses Israel-Palestine issues and appears among the most commented accounts in both the US and Israel splits.
Figure S16: Example of @elonmusk post classified as United Kingdom per the methodology outlined in C.5.1.
Appendix D Outcomes of the algorithmic resolution
In this section we turn our attention to the analysis of notes according to the status they are attributed by X’s Community Notes algorithm, and publicly disclosed by X.
D.1 Outcomes by country
Leveraging the country-mapping of Community Notes described in Section C.5.1, we present in Figure S17 the fraction of posts ending up with a note with Helpful Status across the 30 countries with at least one thousand annotated posts; computation restricted to the notes considered by X’s Community Notes algorithm i.e. after preprocessing  (?).
Figure S17: Fraction of annotated posts with a note with Helpful Status per country; limited to countries with at least a thousand annotated posts, as per C.5.1, after pre-processing and excluding “Note Not Needed” notes.
We observe significant variation between countries, from 29.0% of posts receiving a note with Helpful Status in Turkey and 19.8% in France to 12.6% in Israel and 11.1% in Argentina.
D.2 Outcomes by topic
To achieve finer granularity than the general topic characterization in Sections C.3 & C.4 and circumvent post-deletion observation bias, we examine word-level differences between Community Notes that achieved Helpful Status and those that did not. We compute chi-square statistics for words of four letters or longer in note content, focusing on the ten countries with the most Community Notes. All keywords were translated into English.
United States
  • Over-Represented words: account, scam, Tesla, fake, steal, video, assets, scams, impersonating, pretext, publicity, scammers, elonmusk, promotion, promotions
  • Under-Represented words: Gaza, political, Donald, Israeli, opinion, quot, context, president, vaccines, evidence, Biden, COVID, Israel, Hamas, Trump
Japan
  • Over-Represented words: earthquake, prediction, forecasting, method, fraud, deception, prize, winning, occurrence, weather, happen, sell, atmosphere, message, time
  • Under-Represented words: deliberation, deputy, reaction, problem, authority, election, evaluation, mild, causality, nature, certification, meeting, relief, party, system
United Kingdom
  • Over-Represented words: shipped, products, scam, unfit, stores, store, generated, purchasing, selling, priced, image, drop, video, caution, exercise
  • Under-Represented words: opinion, violence, political, Israel, gender, male, change, quot, trans, government, people, women, Hamas, party, Labour
Spain
  • Over-Represented words: video, image, corresponds, original, scam, photo, TikTok, digitally, cryptocurrencies, famous, Instagram, Biden, altered, artificial, real
  • Under-Represented words: Montero, parties, Hamas, right, Popular, Quiles, convicted, Pedro, Vito, hate, extreme, crime, Ayuso, PSOE, government
Brazil
  • Over-Represented words: bets, guidelines, house, stake, hidden, site, adult, game, tiger, prohibited, games, pirate, gang, violating, report
  • Under-Represented words: tax, Jair, facts, note, freedom, Hamas, left, Israel, minister, federal, deputy, president, Lula, Bolsonaro, government
France
  • Over-Represented words: acts, account, validate, site, thread, steering, photo, user, usurps, scam, date, original, image, blue, redone
  • Under-Represented words: victims, racism, opinion, quot, hostages, right, racist, Hassan, Palestinians, terrorist, Gaza, political, party, left, Hamas
Germany
  • Over-Represented words: celebrities, scam, advertisement, crypto, replicated, advertise, portals, paid, real, fraud, video, page, acts, abused, world
  • Under-Represented words: study, insult, electricity, climate, government, percent, politics, influence, context, mister, Germany, power, Correctiv, party, Habeck
India
  • Over-Represented words: video, account, websites, tokens, original, pump, dump, impersonating, garbage, forex, sell, digitally, schemes, heavy, advance
  • Under-Represented words: party, constitution, spreading, propaganda, opposition, Muslim, political, Modi, Hindus, Rahul, people, government, Hindu, Gandhi, Congress
Argentina
  • Over-Represented words: Instagram, image, photo, video, original, corresponds, Infobae, Franco, digitally, window, Colapinto, patio, altered, false, Tesla
  • Under-Represented words: Javier, measures, system, president, Massa, party, change, quot, senators, debt, candidate, advances, freedom, government, Milei
Israel
  • Over-Represented words: refused, Gvir, return, convicted, police, examples, millions, Avishi, trust, border, Rothman, drowning, protester, execution, raised
  • Under-Represented words: discussed, Yinon, video, writer, spokesman, Jews, mistaken, probably, brothers, quot, actually, Gaza, original, that, instead
We observe that across countries, scam and doctored media-related keywords are over-represented in notes with Helpful Status compared to those that do not, while politically divisive issues, such as the Israel-Palestine conflict, are under-represented in notes with Helpful Status.
To quantify this observation we curated the following sets of keywords (some are intentionally shortened for expanded matching):
  • US Politics: “Biden”,“Kamala”,“Trump”, “Democrat, “Republican”
  • Israel-Palestine Conflict: “Israel”,“Gaza”,“Netanyahu”, “Hamas, “IDF”, “Hezbollah”, “Tel Aviv”, “Palestin”, “UNRWA”, “Nakba”
  • Russia-Ukraine Conflict: “Russi”,“Ukrain”, “Putin, “Zelensky”, “Moscow”, “Kyiv”,“Kremlin”, “NATO”,
  • Scam: “scam”,“fraud”,“impersonating”
These sets of manually curated keywords are not comprehensive and prioritized precision over recall. We then extracted, in the United States, Community Notes containing one of these keywords, and determine for each topic the fraction of notes reaching Helpful Status:
  • US Politics: 6.0% of notes reached Helpful Status.
  • Israel-Palestine Conflict: 5.5% of notes reached Helpful Status.
  • Russia-Ukraine Conflict: 11.7% of notes reached Helpful Status.
  • Scam: 39.0% of notes reached Helpful Status.
The rate at which notes reach Helpful Status is statistically lower for those related to US Politics, Israel-Palestine Conflict, and Russia-Ukraine Conflict compared to those associated with Scam, as assessed by pairwise permutation test, $p<0.005$.
D.3 Outcomes by account
For each X account, we contrast the observed number of posts receiving note reaching Helpful Status with the expected frequency derived from the population average. Using the chi-square statistic, we identify the 10 most significantly over-moderated and under-moderated accounts across five countries: the United States, Japan, Spain, France, and Brazil.
Importantly, these results are conditioned on posts being available for collection. As such, under-moderated accounts may reflect either lower consensus among raters, a higher tendency of users to delete posts deemed misleading (i.e., those associated with notes with Helpful Status), or some combination of both factors.
United States
  • Under-Moderated: @elonmusk (2.6%), @KamalaHQ (0.1%), @VigilantFox (1.6%), @libsoftiktok (3.3%), @EndWokeness (4.0%), @RadioGenoa (2.1%), @KamalaHarris (1.2%), @JoeBiden (1.6%), @POTUS (2.4%), @krassenstein (1.8%)
  • Over-Moderated: @BGatesIsaPyscho (33.1%),
    @iluminatibot (26.1%), @thehealthb0t (28.8%), @DailyLoud (41.4%), @Evony_TKR (78.1%), @DramaAlert (33.1%), @QTHESTORMM (52.0%), @realstewpeters (29.3%), @dom_lucre (23.0%), @tassagency_en (30.1%)
Japan
  • Under-Moderated: @takavet1 (0.0%), @RyuichiYoneyama (0.0%), @tamakiyuichiro (1.9%), @mas__yamazaki (0.0%), @umekichkun (2.0%), @siroiwannko1 (0.0%), @himasoraakane (0.0%), @nekoruck (0.0%), @bad_texter (2.2%), @noruzo_tetsudo (3.4%)
  • Over-Moderated: @kirinjisinken (81.6%), @tweetsoku1 (40.0%), @geoscience16 (47.5%), @kkkfff1234k (45.1%), @hibakuyada (41.3%), @takui_aka (36.2%), @jhmdrei (30.8%), @giantearthquake (39.3%),
    @foodloss_sos (83.3%), @himuro398 (26.4%)
Spain
  • Under-Moderated: @wallstwolverine (2.4%), @eldiarioes (3.4%), @PSOE (3.3%), @IdiazAyuso (0.0%), @sanchezcastejon (2.0%), @iescolar (2.6%), @Santi_ABASCAL (0.0%), @oscar_puente_ (1.3%), @vox_es (2.9%), @AlanBarrosoA (1.4%)
  • Over-Moderated: @ceciarmy (55.0%), @A3Noticias (29.2%), @Tebasjavier (53.3%), @larazon_es (29.8%), @isaacfouto (43.8%), @MediterraneoDGT (26.8%), @abc_es (18.5%), @EFEnoticias (19.0%), @20m (27.3%), @_anapastor_ (27.6%)
France
  • Under-Moderated: @CerfiaFR (5.4%), @RimaHas (3.9%), @AlertesInfos (5.9%), @julienbahloul (0.0%), @CaronAymericoff (3.7%), @EmmanuelMacron (1.5%), @CordierAlice2 (0.0%), @Pediavenir (0.0%),
    @ArnaultRaphael (4.9%), @emma_ducros (0.0%)
  • Over-Moderated: @silvano_trotta (26.6%), @55Bellechasse (60.9%), @VictorSinclair3 (34.7%), @_IDVL (54.5%), @CestTerrifiant (55.6%), @greenpeacefr (66.7%), @BPartisans (32.4%), @Reporterre (57.9%), @75secondes (48.3%), @pierre_jacquel2 (50.0%)
Brazil
  • Under-Moderated: @LulaOficial (0.8%), @GloboNews (3.9%), @vinicios_betiol (1.7%), @pesquisas_elige (1.4%), @felipeneto (2.7%), @revistaforum (0.0%), @jairbolsonaro (0.9%), @PedroRonchi2 (2.9%), @Estadao (3.7%), @gleisi (4.2%)
  • Over-Moderated: @CR7Brasil (76.7%), @7gamesbet (88.9%), @quero_escolher (86.7%), @choquei (21.4%), @AndreGA_Pe (40.7%), @samengo_ofc (72.7%), @siteptbr (26.3%), @updatecharts (40.5%), @joaquimrib1801 (44.8%), @mspbra (27.1%)
Across different countries, a consistent observation emerges: politically divisive accounts —whether associated with
@KamalaHarris, @tamakiyuichiro, @sanchezcastejon, @EmmanuelMacron, or @LulaOficial— display smaller fraction of post with note having reached Helpful Status compared to the general population. Conversely, entertainment accounts (@DramaAlert, @CR7Brasil) and conspiracy theory promoters (e.g., @BGatesIsaPsycho, @silvano_trotta) tend to generate high agreement and exhibit elevated moderation rates.
Appendix E Ideology scaling of X users
To explore how the latent ideology dimension computed by the Community Notes model relates to Left-Right and Anti-Elite positions, we employ the ideological scaling of X users developed by Ramaciotti et al.  (?). This approach extends Barberá’s  (?) methodology to multidimensional settings, enabling identification and calibration with political survey data across countries.
E.1 Multidimensional ideology scaling
We collected, through the API of X, the set of users following Members of Parliament (MPs) on X across 23 countries: France, Germany, United States, United Kingdom, Spain, Brazil, Australia, Italy, Israel, Argentina, Denmark, Sweden, Canada, Poland, the Netherlands, Switzerland, South Africa, Nigeria, Finland, Mexico, New Zealand, Japan, and Belgium. This sampling strategy enables the selection of a large number of users engaged with national politics; we report in Table S1 the number of such users per country. In the following, we consider only users following at least three MPs and that were followed by at least 25 accounts, following Barberá’s protocol in his 2015 studies  (?, ?, ?).
We then estimate the ideal positions of users and MPs by performing, as Barberá  (?), a Correspondence Analysis  (?) on the follower-MP adjacency matrix. As demonstrated by Carroll et al.  (?), this procedure yields approximations to the ideal points estimated through the following homophilic data generation process:
$P\left(\text{U}_{i}\!\rightarrow\!\text{MP}_{j}|\alpha_{i},\beta_{j},\gamma,\varphi_{i},\varphi_{j}\right)=\text{logistic}\!\left(\alpha_{i}\!+\!\beta_{j}-\gamma\|\varphi_{i}-\varphi_{j}\|^{2}\right),$ (S1)
where, $U_{i}\rightarrow MP_{j}$ stands for the event in which user $\text{U}_{i}$ follows $\text{MP}_{j}$, $\alpha_{i}$ is the level of activity of user $\text{U}_{i}$ measured in number of followed accounts, $\beta_{j}$ is the level of popularity of $\text{MP}_{j}$, measured in number of followers, $\gamma$ is a scale variable, and $\varphi_{i}$ and $\varphi_{j}$ are the multidimensional positions of $\text{U}_{i}$ and $\text{MP}_{j}$ in some unobservable multidimensional space. This data generation process is homophilic in the sense that $P\left(\text{U}_{i}\rightarrow\text{MP}_{j}\right)$ increases as the distance between unobservable positions $\|\varphi_{i}-\varphi_{j}\|$ in some latent space decreases.
Most applications of this scaling method, especially those using social media data, rely on a single-dimensional model for unobservable positions $\varphi$, relying the assumption that the the recovered dimension will match the leading axis of polarization in the national setting under study. While this assumption is usually validated in studies set in the United States  (?), empirical spatial analyses have shown that more general settings, in particular those in European countries  (?), cannot be reduced to a single ideological dimension (e.g., Left-Right).
E.2 Calibration and validation
Ramaciotti et al.  (?) extended this scaling procedure to capture multiple political dimensions by considering the subsequent principal directions $\delta_{2},\delta_{3},...$ identified through the Correspondence Analysis. We then calibrate these latent dimensions with political survey data. This step is required because, in Eq. (S1), the probability $P\left(\text{U}_{i}\rightarrow\text{MP}_{j}\right)$ suffers from an identification issue. Indeed, the model is invariant to isometries, such as flips, translations, and rotations on positions $\varphi_{i}$ and $\varphi_{j}$, meaning that any such transformation would yield identical probability values. Consequently, the raw positions $\varphi_{i}$ lack intrinsic political meaning and cannot serve as reliable indicators for specific political stances or ideologies (ideology being defined as an organized structure of attitudes  (?, ?)). For instance, the relation $\varphi_{i}>\varphi_{j}$ could equally indicate that user $i$ is more Left-leaning than user $j$, or precisely the opposite, rendering the model’s output ambiguous without external anchoring. This identification issue is particularly limiting in cross-country comparative analyses where a dimension of interest for comparison (e.g., a Left-Right) dimension, might present itself as different spatial directions in the latent homophily space of positions $\varphi$ in equation Eq. (S1). This limitation is made all the more salient when the phenomenon of interest, such as factuality of claims, is expected to depend on a leading dimension of polarization, and dimensions linked to trust in elites and institutions  (?).
To resolve this identification problem, we use the Global Party Survey  (?)—an expert survey positioning parties along several issue and ideology dimensions—to map positions $\varphi_{j}$ from the latent homophily spaces to positions along the dimensions of the expert survey. Specifically, we map the average position in the latent homophily space of MPs from a party to the position of the party along the dimensions assessed in the Survey.
We consider the Global Party Survey for our study because of its international coverage and available coherence metrics with other studies on party positions. We retain for our study the Left-Right dimension (variable V4_Scale) and the dimension measuring populist rhetoric, capturing attitudes towards elites and institutions (among other populist traits; variable V8_Scale).
To validate the positions of users in our study along the Left-Right position, we leverage an independent signal not used to infer users’ leaning, namely, the textual content of users’ profile bios. Specifically, we prompted a Large Language Model, in our case Zephyr 7B $\beta$  (?), to identify a subset of users that could be attributed with the labels “Left” or “Right”. We display in Figure S18 the distribution of inferred Left-Right leaning for users labeled by the LLM as “Left” or “Right”. Our validation method shows that this Left-Right dimension separates users that can be labeled as binary being either “Left” or “Right” leaning on the basis of what they write in their profile bios. We quantify this spatial separation with the F1 performance of an logistic regression classifier.
Figure S18: Distribution of the inferred Left-Right leaning segmenting users labeled by LLM Zephyr 7B $\beta$  (?) as “Right-leaning” or “Left-leaning”. We report the optimal threshold best segmenting these populations as well as the associated F1-score.
It is noteworthy that, while the Left-Right dimension correctly separates users including Left or Right identification cues in the social media profiles, the order is inverted for Poland in Figure S18. This inversion is anticipated by the literature in political sciences, and observed empirically in different studies. In most Eastern-European post-communist countries, the most structuring political dimension can be identified as spanning from Left-Authoritarian to Right-Libertarian positions, in contrast with Western-European political space structured between Left-Libertarian to Right-Authoritarian positions  (?, ?, ?). For instance, the Law and Justice Party (Prawo i Sprawiedliwość, or PiS), is recognized to be a Right-wing party that supports nonetheless several economically Left-leaning policies  (?). Figure S19 provides further explanation into this inversion in our data. In the survey used for the identification of the latent space of the ideology scaling model (GPS), positions of available parties (three) along Left-Right and Liberal-Conservative are negatively correlated. This means that, in our ideology scaling data of X users, Left-Right and Liberal-Conservative dimensions are related by a flip and translation operation. Among the dimensions available in the GPS data, the Liberal-Conservative dimensions provides the second viable dimension structuring Left or Right identification. An examination of substance of the questions presented to respondents of the survey surface a predominance of economic issues in Left-Right positioning (including, e.g., issues such as taxation, privatization and welfare state), and a predominance of cultural issues in Liberal Conservative positioning (including, e.g., mentions to order, tradition, abortion, and same-sex marriage). Figure S19 further shows that order in self-identification into Left and Right groups is inverted (with respect to other countries) along the Left-Right dimension, while being coherently ordered in the Liberal-Conservative dimension. Because of the sign invariance of the latent ideology dimension in the Community Notes modes (i.e., $\theta_{n}$ and $\theta_{r}$), this inversion has no consequences in our conclusions, as they hinge on the alignment and not the orientation of dimensions.
Figure S19: Additional details for the validation of ideology scaling data for Poland in Figure S19. Left panel shows that, for parties available in the survey used for space identification (Global Party Survey), Left-Right and Liberal-Conservative positions are inverted. Center and right panels show cues of Left and Right identification provided by users in the social media profiles are coherent with the Liberal-Conservative dimension, and inverted for the Left-Right dimension.
E.3 Community Notes population with inferred ideology
Community Notes contributors are not publicly linked to their X handles, preventing direct comparison between rater latent ideology $\theta_{r}$ and their political positions inferred from followed MPs. Instead, we compare the latent ideology of each note $\theta_{n}$ with the ideology of the X account that authored the original post to which the note is attached.
To enable an analysis disaggregated by national setting, we segment posts by country. Specifically, we define a post and its associated Community Note as involved in a country’s political debate when the author of the post follows at least three Members of Parliament from that country and the post is written in one of the country’s national languages.
This country segmentation methodology differs from the approach outlined in Section C.5.1, serving distinct purposes and presenting different limitations. This procedure assigns country classifications at the post level, which then extends to all associated notes, necessarily requiring post metadata. This creates a practical constraint: as of March 2025, 18.2% of unique posts with Community Notes were unavailable for collection —due to their deletion— preventing author identification and country assignment. As detailed previously in Section B.2, posts associated with note with Helpful Status are more likely to be deleted. This limitation explains our preference for note-specific country mapping whenever possible, particularly when characterizing global Community Notes usage patterns. Nevertheless, by necessity we employ post-specific country mapping to enable comparisons between X’s learned ideological representations and established ideological scaling datasets.
Among the 23 geographically diverse countries initially considered in this study, listed in Section E.1, we focus on the 13 countries that have at least one thousand unique X accounts receiving Community Notes with inferred ideology: United States, United Kingdom, Japan, Spain, France, Brazil, Canada, Germany, Argentina, Israel, Australia, Poland, and Mexico. We report in Table S1 the number of notes, posts, and X accounts with inferred ideological positions across all 13 countries.
Country Number of Number of Number of Posts Total Number of Accounts
Notes Posts Authors with Inferred Ideology
Argentina 12578 9430 1819 1410839
Australia 11927 9064 1345 554224
Brazil 28690 21188 2925 6142800
Canada 25797 18991 2603 786589
France 32152 25184 3763 979089
Germany 25712 19663 3298 1169464
Israel 14920 11194 1686 407378
Japan 27397 21960 4259 1489166
Mexico 7500 5643 1275 2120278
Poland 10928 8704 1536 732669
Spain 38289 28377 4095 1723836
Uk 59352 43655 6813 2007915
Us 166951 117235 14582 10927393
Table S1: Number of notes, posts, post authors, and the total number of X accounts with inferred ideological positions (including those without Community Notes) by country.
Figure S20: Distribution of the inferred Left-Right leaning of X users who received at least one proposed Community Note. We report the average Left-Right leaning of the annotated population (plain grey line) and of the global population following 3 MPs (dotted red line).
Figure S20 displays the distribution of Left-Right ideology among X users associated with at least one Community Note (irrespective of “helpfulness” status), overlaid with the average Left-Right leaning of the general population following MPs. This comparison shows general alignment between users associated with proposed Community Notes and the broader population of users following at least three national MPs. Additionally, across all 13 countries in our dataset, we report in Figure S21 the distribution of Left-Right ideology among X accounts having authored posts for which Community Notes i) were requested, ii) were proposed, and iii) reached Helpful status.
Figure S21: Distribution of the inferred Left-Right ideology of X users having authored posts for which Community Notes i) were requested, ii) were proposed, and iii) reached Helpful status.
Appendix F Inferring the Community Notes model
We now turn to the analysis of the parameters inferred by X’s Community Notes algorithm from contributor ratings.
F.1 Matrix factorization
X’s Community Notes system aims to surface notes deemed Helpful by users across “diverse viewpoints”  (?), implementing a bridging-based ranking approach  (?). This objective is operationalized through matrix factorization, which learns an opinion space, positioning raters and notes in an unsupervised manner, to explain observed ratings. The core matrix factorization underlying X’s Community Notes system is formalized in Equation S2  (?):
$\underbrace{\hat{\eta}_{rn}}_{\text{Predicted Rating}}=\underbrace{\beta_{0}}_{\text{Global Bias}}+\underbrace{\beta_{r}}_{\begin{subarray}{c}\text{Rater Bias}\\ \text{``leniency"}\end{subarray}}+\underbrace{\beta_{n}}_{\begin{subarray}{c}\text{Note Bias}\\ \text{``helpfulness"}\end{subarray}}+\underbrace{\theta_{n}\cdot\theta_{r}}_{\begin{subarray}{c}\text{Note-Rater}\\ \text{Ideological Alignment}\end{subarray}}$ (S2)
With the actual ratings $\eta_{un}$ numerically encoded as 1 for Helpful, 0.5 for Somewhat Helpful and 0 for Not Helpful. The baseline ratings are captured by the bias terms, $\beta_{0},~{}\beta_{r},~{}\beta_{n}$. A high value of $\beta_{r}$ indicates an high propensity for the rater $r$ to rate notes as Helpful, while a high $\beta_{n}$ reflects a note’s tendency to receive Helpful ratings across the rater base. $\theta_{n}$ and $\theta_{r}$ encode notes and raters latent ideology.
The product $\theta_{r}\cdot\theta_{n}$ quantifies ideology alignment, with the rational that when raters share the note’s ideology, they are more likely to rate it as Helpful. When $|\theta_{n}|$ is large, voting behaviors on that note varies a lot in function of raters’ latent ideology $\theta_{r}$. To maximize this effect, X applies, during training, stronger regularization on bias terms $\beta_{0},~{}\beta_{r},~{}\beta_{n}$ compared to latent ideology terms $\theta_{r},~{}\theta_{n}$  (?, ?).
Through this design, when ratings on a note depend on raters’ latent ideology (e.g., raters with $\theta_{r}>0$ rate the note Helpful while those with $\theta_{r}<0$ rate it Not Helpful), the note latent ideology $\theta_{n}$ will align with raters’ (e.g. $\theta_{n}>0$). Correspondingly, the bias term $\beta_{n}$ will be small due to regularization.
Conversely, when consensus emerges (i.e. raters with different $\theta_{r}$ signs agree), $|\theta_{n}|$ will shrink to minimize rater leaning impact on predicted ratings $\hat{\eta}_{rn}$, while $|\beta_{n}|$ will be high: if raters agree the note is Helpful then $\beta_{n}>0$, if they agree it is Not Helpful then $\beta_{n}<0$.
Finally, X’s Community Notes system assign Helpful Status/Not Helpful Status to notes when their bias term $\beta_{n}$ exceeds specified thresholds. For intermediate cases where rater leaning $\theta_{r}$ effectively predicts rating patterns, X’s system withholds classification pending additional diverse ratings, so-called Needs More Ratings Status.
F.2 Computational implementation
While X’s Community Notes algorithmic implementation is open-source  (?), it incorporates multiple heuristics and experimental features that increase model complexity. For simplicity, we solely implement the core matrix factorization from Equation S2. We will validate inferred values of the Community Notes models by comparing the outcomes produced by the model we train, with those observed in the data disclosed by X.
Following X’s approach  (?, ?), we learn biases $\beta_{0},~{}\beta_{r},~{}\beta_{n}$ and latent ideologies $\theta_{r},~{}\theta_{n}$ by minimizing the loss function in Equation S3, with regularization five times higher on biases than on latent ideologies:
$\min_{\{\beta,\theta\}}\sum_{\eta_{rn}}\underbrace{(\eta_{rn}-\hat{\eta}_{rn})^{2}}_{\text{Reconstruction Error}}+\quad\underbrace{5\lambda(\beta_{0}^{2}+\beta_{r}^{2}+\beta_{n}^{2})+\lambda(\theta_{r}^{2}+\theta_{n}^{2})}_{\text{Regularization}}.$ (S3)
In their publication, X’s researchers justify higher regularization on biases compared to latent ideologies to minimize “risks to [the] community and reputation from increasing visibility of low quality notes”  (?). Indeed, by increasing regularization on note and rater biases ($\beta_{n}$, $\beta_{r}$), this encourages latent ideologies ($\theta_{n}$, $\theta_{r}$) to explain more of the rating patterns. As such, the bias terms, particularly $\beta_{n}$, can be interpreted as the idiosyncratic “quality” of the note—-i.e., the propensity of raters to rate it “helpful” beyond their ideological preferences. However, we should emphasize that the numerical factor of 5 in Equation (S3) is arbitrarily set, and tweaking it can directly impact the fraction of notes perceived to “foster consensus”.
Contrary to X’s production system that processes ratings sequentially, we aggregate all ratings until March 1, 2025. This larger training dataset requires adjusting the regularization magnitude $\lambda$ compared to X’s production system.
Figure S22: Hyperparameter tuning of the regularization strength $\lambda$ (A) and number of training epoch (B) evaluated using reconstruction error on a held-out dataset.
The validation loss plateaus quickly during the training, to ensure convergence and model stability, we performed three epochs. We tuned the hyperparameter $\lambda$, while preserving X’s 5:1 ratio between biases and latent ideologies regularization, by measuring the reconstruction error $|\eta_{rn}-\hat{\eta}_{rn}|$ on 90/10 train/test splits, after three epochs. In Figure S22, the optimal value of $\lambda=2.5\times 10^{-5}$ is identified.
Informed by the hyperparameter tuning, we then trained the model for three epochs with a regularization strength of $\lambda=2.5\times 10^{-5}$ at a learning rate of $2.5\times 10^{-3}$. After training, the average reconstruction error $|\eta_{un}-\hat{\eta}_{un}|$ between the predicted rating $\hat{\eta}_{rn}=\beta_{0}+\beta_{n}+\beta_{r}+\theta_{n}\cdot\theta_{r}$ and raters’ effective ratings $\eta_{un}$ averages at $0.204$. Figure S23 displays the average reconstruction error per country (segmenting according to the methodology in C.5.1), showcasing comparable performance across countries.
Figure S23: Average reconstruction error $|\eta_{un}-\hat{\eta}_{un}|$, between the predicted rating and raters’ effective ratings, per country (see S.C.5.1).
F.3 Validating the parameters of the model
The distribution of notes and rater biases and latent ideology inferred from Community Notes ratings, displayed in Figure S25, mirrors those reported in X’s article  (?) and aligns with expectations from X’s Community Notes source code  (?).
Figure S24: Distribution of notes (raters) bias $\beta_{n}$ ($\beta_{r}$) and latent ideology $\theta_{n}$ ($\theta_{r}$).
We segment in Figure S25 notes by X’s Community Notes outcomes: Helpful Status, Needs More Ratings Status, or Not Helpful Status. Conform to expectation notes with Helpful Status have high positive bias $\beta_{n}$, those with Not Helpful Status have large negative bias, and notes with Needs More Ratings Status have high latent ideology $|\theta_{n}|$. Quantitatively, 90% of notes with Helpful Status have bias terms $\beta_{n}>0.180$, 90% of notes with Not Helpful Status have $\theta_{n}<-0.159$, and 69% of notes with Needs More Ratings Status fall between these thresholds. The inferred note bias, $\beta_{n}$ allows prediction of whether X assigned a note Helpful Status with One-vs-the-rest AUC of 0.92 and Not Helpful Status with One-vs-the-rest AUC of 0.97.
Figure S25: Distribution of notes bias $\beta_{n}$ and latent ideology $\theta_{n}$, segmented by status assigned by X: Helpful Status, Needs More Ratings Status, or Not Helpful Status.
The differences between the numerical thresholds and those in  (?) stem from the difference in dataset size. While we learn note and rater representations across all available ratings, X’s Community Notes production system aims to surface, in a timely fashion, Helpful notes to users by considering recent ratings. Additionally, X implements heuristics constraining note status changes, which we did not implement, focusing solely on the core matrix factorization of Equation S2.
Appendix G Analysis of the latent ideology dimension learned by Community Notes
The matrix factorization used by X’s Community Notes system aims to learn a continuous opinion space from rating patterns  (?). Given that Community Notes predominantly address political content (see Section C.3), this learned opinion space is expected to capture raters’ and notes’ political leanings and the surfaced notes to be rated as Helpful by users across the political spectrum. Through surveys and self-reported political affiliations, X’s researchers validated this objective during the pilot project  (?). Allen et al.  (?) further substantiated —during the pilot phase with $\sim 5000$ notes— these findings by showing that users political partisanship strongly predicts their ratings.
Despite growing literature on crowd-sourced fact-checking, the opinion space learned by X’s Community Notes system —which drives note selection— remains unexamined. This section characterizes this learned space and evaluates its alignment with Twitter users’ political leanings. To this end, we first present in Section E.1 the ideological scaling data leveraged in this analysis, and then compare it with the learned representations in Sections G.1 & G.2.
G.1 Left-Right dimension
To foster participation in Community Notes, contributors’ profiles are not publicly linked to their X accounts, preventing direct comparison between rater latent ideology $\theta_{r}$ and their political positions inferred from followed MPs. Consequently, we are left to compare the latent ideology of Community Notes $\theta_{n}$ with the political leaning of X accounts that authored the posts to which notes are attached, restricting our analysis to non-deleted posts.
As a first qualitative analysis, we report the position of popular accounts in five countries over this latent ideology scale. Specifically, for each country, we compute the average latent ideology $\theta_{n}$ of community notes associated with each X account’s posts. We then divide the range of average $\theta_{n}$ values into 10 bins and manually inspect accounts within each bin. To protect privacy while providing illustrative examples, we focus on accounts that have received at least 10 community notes (irrespective of status), have the highest follower counts within their respective bins, and whose political leanings are manifestly public. To validate the reliability of the average latent ideology $\theta_{n}$, one can observe that for accounts with at least 10 notes, 83.6% of those notes have the same latent ideological sign ($\text{sign}(\theta_{n})$).
Figure S26: Average positions of notes added to posts of political figures and news outlets per country. The position of accounts depend on the average latent ideology of notes $\theta_{n}$ in the Community Notes models in the United States, Brazil, Japan, France and Spain, given to them.
We observe in Figure S26 that Community Notes associated with posts published by Right-leaning accounts tend to have negative latent ideology ($\theta_{n}<0$), while those associated with Left-leaning authored posts have positive latent ideology ($\theta_{n}>0$). For instance, notes added to posts by U.S. Vice President J. D. Vance, Brazil’s former President Jair Bolsonaro, France’s far-Right leader Jordan Bardella, and Spain’s far-Right party Vox have negative latent ideology $\theta_{n}$. Driven by homophily, this indicates that Left-leaning Community Notes contributors have negative latent ideology $\theta_{r}<0$, so that they rate Helpful notes under Right-leaning posts but Not Helpful notes under Left-leaning posts, and conversely Right-leaning contributors have positive latent ideology $\theta_{r}>0$.
The second observation is that, while different countries could in principle exhibit different sign orientations in how $\theta_{r}$ and $\theta_{n}$ map onto Left-Right, all countries in our study share the same orientation: $\theta_{n}<0$ under post of Right-leaning users and $\theta_{n}>0$ under post of Left-leaning users. Importantly, this alignment emerges naturally from the non-supervised inference procedure without any explicit constraints. This observation can be explained by cross-country rating patterns. Since the Community Notes algorithm considers ratings as a whole rather than segmenting them by country, raters’ latent ideological positions $\theta_{r}$ must align across national settings to best explain their rating behavior.
To corroborate this interpretation, we display cross-country rating patterns in Figure S27. Specifically, we show the fraction of users who rated at least one note from a given country and also rated at least one note from another country. We observe that an overwhelming proportion of contributors rated US notes in addition to notes from their own countries.
Figure S27: Cross-country rating patterns among Community Notes contributors. The figure shows the percentage of users who rated at least one note from a given country (columns) and also rated at least one note from another country (rows). For instance, 59.6% of users who rated at least one Japanese Community Note also rated at least one US note, while only 7.3% of users who rated at least one US note also rated at least one Japanese note.
It should be emphasized that the predicted ratings $\hat{\eta}_{rn}$ in Equation S2 are invariant to a joint sign inversion of latent ideologies: $(\theta_{n},\theta_{r})\mapsto(-\theta_{n},-\theta_{r})$. As such, while we report Left-leaning contributors as having negative latent ideology ($\theta_{r}<0$) and Right-leaning contributors as having positive latent ideology ($\theta_{r}>0$), the exact opposite assignment could hold without any impact on the model’s operation. All subsequent discussion holds without loss of generality under either sign convention.
G.1.1 United States
Alignment with the Left-Right dimension in political survey data.
In the United States, 95.0% of Community Notes added to posts authored by Republican members of 118th United States Congress have negative latent ideology $\theta_{n}<0$, while 94.6% of those associated with Democratic members of Congress have positive latent ideology $\theta_{n}>0$ (excluding “Note Not Needed” notes).
Similarly, considering accounts following at least 10 US MPs, 82.2% of notes given to posts published by accounts following a majority of Democratic MPs have a positive latent ideology ($\theta_{n}>0$); while 85.1% of notes given to posts published by accounts following a majority of Republican MPs have a negative latent ideology ($\theta_{n}<0$).
Relying on the continuous Left-Right dimension inferred in Section E we can further quantify this alignment. When framed as a prediction task, the position of a user on the Left-Right political spectrum in the United States predicts the sign of $\theta_{n}$ (the latent ideology of the Community Notes associated with their post) with an AUC-ROC of $0.822\pm 0.009$ (standard deviations estimated through 10-fold cross-validation).
To corroborate the Left-Right alignment of note latent ideology, we examine a signal not leveraged in the inference of such ideology: note authorship. Indeed, X Community Notes system only considers ratings, not which contributor wrote the community notes being rated. Therefore, examining authorship pattern of contributors with $\theta_{r}>0$ or $\theta_{r}<0$ allows us to further characterize the underlying dynamics. We observe in Figure S28 that users with $\theta_{r}<0$, who rate Helpful notes under Right-leaning accounts, are also the ones writing notes under Right-leaning accounts; aligning with the observations made by Allen et al. during the pilot program  (?). Conversely, users with $\theta_{r}>0$, who rate Helpful notes under Left-leaning accounts, are also the ones writing notes under Left-leaning accounts.
Figure S28: Distribution of Left-Right leaning of X accounts having authored posts to which Community Notes were written by contributors with $\theta_{r}<0$ (plain blue) and $\theta_{r}>0$ (dashed red).
Media Bias.
The alignment between the ideological space learned by X’s Community Notes system and established political dimensions can be further substantiated relying on a signal independent from poster following patterns, namely, the use of news media outlet as sources in Community Notes.
Specifically, we consider the 109 US news media outlets, used in at least 100 Community Notes, whose ideological skew was also evaluated by Ad Fontes Media  (?). We subsequently compare the Left-Right bias of the news outlets, as assessed by Ad Fontes Media, and the average latent ideology $\theta_{r}$ of the Community Notes contributors who authored notes using these media as sources. As displayed in Figure S29, we observe a strong Pearson correlation between these quantities: $\rho=0.744$ $p<10^{-4}$, with a $[0.65,~{}0.82]$ Confidence Interval at 95% for correlation, estimated via Fisher transformation.
Importantly, this analysis relies on sources used in Community Notes —data disclosed by X— without requiring additional data collection compared to post-based metrics. This approach avoids biases from users removing posts after the note reached Helpful Status notes, providing complementary robustness to the previously documented alignment.
Figure S29: Ad Fontes Media  (?) Left-Right bias for news media outlets as function of the average latent ideology $\theta_{r}$ of Community Note contributors who authored notes using those domains as sources. Negative values indicate Left-leaning bias, positive values indicate Right-leaning bias. The size of each node is proportional to the number of Community Notes referencing each domain.
Fact-Checks.
We inspect the latent ideology distribution of Community Notes referencing expert fact-checking authored by 160 organizations; see curation in Section C.2.2. Figure S30 displays the distribution of note latent ideology $\theta_{n}$ and of contributors who authored the Community Notes $\theta_{r}$. We observe that Community Notes containing expert fact-checks span the full ideological axis. Yet, a majority of notes are authored by Left-leaning contributors, specifically $60.3\%~{}[58.0,~{}62.8]$ of notes are authored by contributors with $\theta_{r}<0$; 95% confidence interval over 100 bootstraps with replacement over the notes and the set of fact-checking organizations considered in the analysis.
Notes incorporating expert fact-checking sources reach Helpful Status at significantly higher rates ($21.3\%$
$[18.3\%,~{}24.1\%]$) than average, with this effect more pronounced for notes authored by Right-leaning contributors ($\theta_{r}>0$): $26.9\%[23.3\%,~{}30.8\%]$ versus $17.6\%~{}[14.0\%,~{}20.7\%]$ for Left-leaning contributors ($\theta_{r}<0$); 95% confidence interval over 100 bootstraps with replacement over the notes and the set of fact-checking organizations considered in the analysis.
Figure S30: Distribution of note latent ideology $\theta_{n}$ (A) and note author latent ideology $\theta_{r}$ (B) for notes containing expert-authored fact-checks.
Deletion Rate.
In the main text, we reported that notes with Helpful Status appeared more frequently on posts published by Right-leaning X accounts compared to Left-leaning accounts. Specifically, in the United States, 8.3% of annotated posts authored by Left-leaning users receive notes that achieve Helpful Status, compared to 12.9% for Right-leaning users, similar to what reported Renault et al.  (?). However, this observation requires careful interpretation due to a critical methodological limitation: we can only infer the political leaning of an account if the X post was available at the time of data collection. When posts are deleted, the identity of the account that authored the post discussed in the note is not disclosed.
The observed asymmetry may be explained, at least partially, by differences in post deletion behavior. If Left-leaning users delete their posts after being flagged as “misinforming or potentially misleading” more frequently than Right-leaning users, an observer would perceive that Left-leaning users are less frequently flagged for misinformation.
We illustrate this bias by constructing a fictitious example that matches our own observations. Consider a scenario where the true rate of posts receiving notes with Helpful Status ($f_{\text{helpful}}$) is 13% for both Left and Right-leaning users, and both delete their non-misleading posts at the same rate ($d_{\text{not helpful}}=15\%$). However, suppose Left-leaning users delete $d^{\text{Left}}_{\text{helpful}}=48.5\%$ of their posts associated with notes having Helpful Status, while Right-leaning users delete only $d^{\text{Right}}_{\text{helpful}}=16\%$ of such posts.
Under these conditions, the observed proportion of notes having Helpful Status associated with Left-leaning accounts would be:
$\text{Observed}^{\text{Left}}_{\text{helpful}}$ $=\frac{f_{\text{helpful}}\times(1-d^{\text{Left}}_{\text{helpful}})}{f_{\text{helpful}}\times(1-d^{\text{Left}}_{\text{helpful}})+(1-f_{\text{helpful}})\times(1-d_{\text{not helpful}})}$
$=\frac{0.13\times(1-0.485)}{0.13\times(1-0.485)+(1-0.13)\times(1-0.15)}$
$=8.3\%$
Similarly, for Right-leaning accounts:
$\text{Observed}^{\text{Right}}_{\text{helpful}}$ $=\frac{0.13\times(1-0.16)}{0.13\times(1-0.16)+(1-0.13)\times(1-0.15)}$
$=12.9\%$
These hypothetical values align with our empirical observations, suggesting that differential deletion rates could explain the apparent asymmetry.
To investigate this possibility, we leveraged the ideological positioning inferred by X’s Community Notes algorithm. All notes, including those associated with (now) deleted posts, are positioned on the same ideological axis, and we previously showed that notes with $\theta_{n}<0$ are generally associated with posts authored by Right-leaning accounts, while notes with $\theta_{n}>0$ are associated with posts authored by Left-leaning accounts.
Segmenting notes based on their inferred ideology $\theta_{n}$ as a proxy for the post authors’ political leaning, we observe distinct deletion patterns:
Posts (presumably written by Right-leaning accounts) associated with Community Notes with $\theta_{n}<0$ :
  • having reached Helpful Status, were deleted in 35.9% of cases
  • having not reached Helpful Status, were deleted in 16.9% of cases
Posts (presumably written by Left-leaning accounts) associated with Community Notes with $\theta_{n}>0$ :
  • having reached Helpful Status, were deleted in 45.2% of cases
  • having not reached Helpful Status, were deleted in 19.3% of cases
A permutation test over note latent ideology signs ($sign(\theta_{n})$) confirmed the statistical significance of these differences.
Under these deletion rates, Left-leaning accounts would appear to have notes with Helpful Status on 8.3% of their retrievable posts if the true rate is 11.8%, while Right-leaning accounts would appear to have such notes on 12.9% of their retrievable posts if the true rate is 15.6%.
Importantly, we cannot extrapolate these deletion rates to the general population. As described by Wojcieszak et al.  (?), “most users do not follow political elites on X; those who do show overwhelming preferences for ideological congruity.” Therefore, Left-leaning accounts heavily invested in political discussions and following multiple MPs (and thus positioned on our Left-Right scale) may delete posts flagged as misleading at rates even higher than the estimate of 45.2%.
Finally, it should be noted that X’s Community Notes algorithm is blind to the identity of both the author of the post and that of the contributor having proposed the note, inferring an ideological space for notes and raters solely based on rating patterns. Because of the mathematical symmetry between ($\theta_{n}$, $\theta_{r}$) and (-$\theta_{n}$, -$\theta_{r}$) in Equation S2, there is no a priori reason to assume different “helpfulness” rates based on latent ideology. In fact, the fraction of Community Notes with $\theta_{n}<0$ reaching Helpful Status does not statistically differ from those with $\theta_{n}>0$, and the same holds for notes authored by contributors with $\theta_{r}<0$ and $\theta_{r}>0$.
Considering the evidence at hand, rather than interpreting the observed differences as definitive evidence that Right-leaning users are flagged more often than Left-leaning users for sharing misinformation, as did Renault et al.  (?), we should carefully understand these differences as composite measures reflecting both the true rates of agreement leading to Helpful Status of notes and differential deletion behaviors across the political spectrum. Based on collected public information, it is not possible to quantify with certainty any potential partisan asymmetry in misinformation sharing on X as flagged by Community Notes. Further experiments collecting posts as they receive community notes, prior to reaching Helpful Status, would allow settling this alternative explanation.
G.2 Anti-Elite dimension
We extend the previous analysis to all 13 countries for which we possess ideological scaling data with significant overlap with submitted Community Notes. We include in our analysis an additional dimension identified in the literature as relevant for the spread of misinformation  (?). Through the AUC-ROC metric, we quantify the predictive power of various political dimensions in determining the sign of a Community Notes latent ideology $\theta_{n}$ based on the position of the post author along three dimensions:
  • $\delta_{1}$, the most structuring ideological dimension and which best explains MPs following patterns;
  • Left-Right dimension of the Global Party Survey;
  • Anti-Elite dimension of the Global Party Survey.
As shown in Table S2, in each country, the latent ideology $\theta_{n}$ learned by X’s Community Notes system closely aligns with the national structuring political dimension $\delta_{1}$, and, to varying extents, with established political dimensions, i.e., Left-Right, and Anti-Elite.
For instance, in France, the ideological dimension learned by X’s Community Notes system aligns strongly with users’ Left-Right views, with an AUC-ROC of $0.789\pm 0.023$, but weakly with users’ stance on Anti-Elite (AUC-ROC of $0.590\pm 0.045$). Conversely, in Argentina, Left-Right leaning poorly predict note latent ideology sign, while an X account’s position on the Anti-Elite axis predicts the latent ideology sign of associated Community Notes with a high AUC-ROC of $0.859\pm 0.016$.
Country AUC $\delta_{1}$ AUC Left-Right AUC AntiElite AUC Left-Right+AntiElite
Argentina 0.863 ± 0.025 0.682 ± 0.031 0.860 ± 0.024 0.863 ± 0.026
Australia 0.825 ± 0.036 0.809 ± 0.039 0.806 ± 0.033 0.824 ± 0.034
Brazil 0.831 ± 0.027 0.726 ± 0.032 0.552 ± 0.029 0.767 ± 0.035
Canada 0.836 ± 0.033 0.839 ± 0.027 0.832 ± 0.025 0.842 ± 0.025
France 0.775 ± 0.020 0.787 ± 0.016 0.589 ± 0.031 0.816 ± 0.019
Germany 0.840 ± 0.019 0.817 ± 0.020 0.772 ± 0.027 0.852 ± 0.021
Israel 0.729 ± 0.037 0.624 ± 0.037 0.719 ± 0.047 0.718 ± 0.041
Japan 0.731 ± 0.024 0.760 ± 0.020 0.715 ± 0.023 0.757 ± 0.020
Mexico 0.818 ± 0.031 0.758 ± 0.035 0.814 ± 0.030 0.822 ± 0.028
Poland 0.850 ± 0.041 0.838 ± 0.039 0.846 ± 0.042 0.850 ± 0.040
Spain 0.842 ± 0.023 0.850 ± 0.018 0.787 ± 0.021 0.855 ± 0.019
UK 0.780 ± 0.015 0.793 ± 0.015 0.735 ± 0.017 0.793 ± 0.015
US 0.822 ± 0.009 0.822 ± 0.008 0.822 ± 0.008 0.822 ± 0.008
Table S2: AUC-ROC in predicting the majority latent ideology sign ($\text{sign}(\theta_{n})$) of notes associated with X accounts based on their position along the $\delta_{1}$ axis, Left-Right axis, Anti-Elite axis, and both Left-Right and Anti-Elite axes combined. Standard deviations estimated through 10-fold cross-validation.
For greater interpretability, we project in Figure S31 the X accounts onto the 2D plane defined by the left-right and Anti-Elite axes, then trained a logistic regression model to predict the sign of the latent ideology of Community Notes associated with X accounts using their coordinates on this plane as input. To bolster the robustness of the analysis, we performed 10-fold cross validation and report the average AUC-ROC of this prediction task with standard deviation across the 10 folds. Overall, with an average AUC-ROC of $0.814\pm 0.034$, we observe that across national political settings, X’s Community Notes system successfully captures users’ ideological leanings.
Figure S31: X accounts positioned in the (Left-Right; Anti-Elite) 2D plane, segmented by the majority sign of the latent ideology $\theta_{n}$ of their associated Community Notes (red for $\theta_{n}<0$ and blue for $\theta_{n}>0$). The direction that best separates notes by majority outcome, learned through logistic regression, is displayed as $CN$ along with the corresponding AUC score for sign prediction based on account positions in the 2D plane (standard deviation from 10-fold cross-validation). The direction structuring each country’s political landscape $\delta_{1}$ is projected onto the plane.
Finally, as coherent validation, we rely on a signal not considered in X Community Notes systems: note authorship. We display in Figure S32 the position along the structuring dimension $\delta_{1}$ of X accounts associated with Community Notes authored by contributors with $\theta_{r}<0$ and $\theta_{r}>0$. Overall, we observe a stark distinction between these two sets of contributors in the UK, Spain, France, Brazil, Canada, Argentina, Poland, and Mexico, while the distinction is less pronounced in Japan, Germany, Israel, and Australia. This indicates, in relation to Table S2, that in the latter countries, the authoring of Community Notes is not as aligned with the rating of Community Notes, or alternatively that ratings of Community Notes are not structured along the same dimension as MP-following patterns.
Figure S32: Distribution of the position of X accounts along $\delta_{1}$, segmented by the Community Notes author latent ideology sign ($\theta_{r}<0$ plain blue and $\theta_{r}>0$ dashed red).